A 220v motor draws current of 2.8A.determine the electrical energy used in 1hours
1
Expert's answer
2020-03-02T10:23:32-0500

The electrical energy used by a motor is equal to work done by an electric current


"W=IVt=2.8\times 220\times 3600=2.22\times 10^6\:\rm J=2.22\:\ MJ."